Question 1 Find (i) ∫ 𝑑π‘₯ π‘₯(ln π‘₯)2, (ii) ∫ π‘₯√π‘₯ + 1 𝑑π‘₯, (iii) ∫ π‘₯(π‘₯ 2 + 3) 4𝑑π‘₯ 1 0 , (iv) cos2 π‘₯ sin3 π‘₯ 𝑑π‘₯. Try 𝑑 = sin x


1
Expert's answer
2021-10-11T06:01:32-0400

"\displaystyle
2\int x\ln x \\
\text{Using integration by parts, we evaluate the above integral}\\
\text{Let $u = \ln x, dv = x$ }\implies du = \frac{1}{x}, v = \frac{x^2}{2}\\
\int udv = uv -\int vdu\\
\implies 2\int x\ln x = 2(\frac{x^2}{2}\ln x - \frac{x^2}{4}+c)\\
x^2\ln x - \frac{x^2}{2}+c\\
\int (x\sqrt{x}+1)dx = \int (x^{\frac{3}{2}}+1)dx\\
= \frac{2x^{\frac{5}{2}}}{5}+x+c\\
\int x(x^2+3)4dx= \int (4x^3 +12x)dx\\
x^4+6x^2+c\\
\int \cos 2x \sin 3x dx\\
\sin A \cos B = \frac{1}{2}(\sin 5x + \sin x)dx\\
\therefore \int \cos 2x \sin 3x dx = \frac{1}{2}\int \sin 5xdx +\frac{1}{2} \int \sin xdx\\
=-\frac{1}{10}\cos5x-\frac{1}{2}\cos x+c"